A superconducting solenoid is wound with 200 turns
Answer #1

here,

number of turns , N = 200

length of wire , L = 0.15 m

magnetic feild , B = 15 T

let the current throug solenoid be i

B = u0 * N*i/L

15 = 1.26 * 10^-6 * 200 * i/0.15

i = 9 * 10^3 A

the current through the solenoid must be (a) 9 * 10^3 A